# LAA-Net: Localized Artifact Attention Network for High-Quality Deepfakes Detection

### v1

@@ Abstract
This paper introduces a novel approach for high-quality deepfake detection called Localized Artifact Attention Network (LAA-Net). Existing methods for high-quality deepfake detection are mainly based on a supervised binary classifier coupled with an implicit attention mechanism. As a result, they do not generalize well to unseen manipulations. To handle this issue, two main contributions are made. First, an explicit attention mechanism within a multi-task learning framework is proposed. By combining heatmap-based and self-consistency attention strategies, LAA-Net is forced to focus on a few small artifact-prone vulnerable regions. Second, an Enhanced Feature Pyramid Network (E-FPN) is proposed as a simple and effective mechanism for spreading discriminative low-level features into the final feature output, with the advantage of limiting redundancy. Experiments performed on several benchmarks show the superiority of our approach in terms of Area Under the Curve (AUC) and Average Precision (AP). The code will be released soon.
@@ C H E-FPN Test set AUC (%)
95.56 98.54 82.21 74.98 87.82
79.87 94.60 71.70 72.47 79.66
91.56 98.27 78.35 73.02 85.30
95.40 98.43 86.94 80.03 90.20
Table 3: Pair-wise ablation study under the cross-dataset setup of the Consistency branch (C), Heatmap branch (H), and multi-layer fusion with E-FPN.
@@
Table 4: Traditional FPN versus E-FPN, using the SBI-based data synthesis under the cross-dataset protocol. Bold and Underline indicate the best and the second best performance, respectively. We report the results when integrating features $\mathbf{F}^{(i)}$ from different layers.
@@ 8 Additional Results
Table 6 and Table 7 report the results under the in-dataset and the cross-dataset seetings, respectively. Overall, it can be seen that LAA-Net achieves better performances than other state-of-the-art methods.

### v2

# LAA-Net: Localized Artifact Attention Network for Quality-Agnostic and Generalizable Deepfake Detection

@@ Abstract
This paper introduces a novel approach for high-quality deepfake detection called Localized Artifact Attention Network (LAA-Net). Existing methods for high-quality deepfake detection are mainly based on a supervised binary classifier coupled with an implicit attention mechanism. As a result, they do not generalize well to unseen manipulations. To handle this issue, two main contributions are made. First, an explicit attention mechanism within a multi-task learning framework is proposed. By combining heatmap-based and self-consistency attention strategies, LAA-Net is forced to focus on a few small artifact-prone vulnerable regions. Second, an Enhanced Feature Pyramid Network (E-FPN) is proposed as a simple and effective mechanism for spreading discriminative low-level features into the final feature output, with the advantage of limiting redundancy. Experiments performed on several benchmarks show the superiority of our approach in terms of Area Under the Curve (AUC) and Average Precision (AP). The code is available at https://github.com/10Ring/LAA-Net.
@@ C H E-FPN Test set AUC (%)
$\times$ $\times$ $\times$ 74.54 92.24 70.81 59.81 74.35
$\times$ $\checkmark$ $\times$ 80.89 94.53 77.93 67.12 80.12($\uparrow$5.77)
$\times$ $\times$ $\checkmark$ 84.21 95.03 80.68 65.47 81.35($\uparrow$7.00)
$\times$ $\checkmark$ $\checkmark$ 95.56 98.54 82.21 74.98 87.82($\uparrow$13.47)
$\checkmark$ $\times$ $\checkmark$ 79.87 94.60 71.70 72.47 79.66($\uparrow$5.31)
$\checkmark$ $\checkmark$ $\times$ 91.56 98.27 78.35 73.02 85.30($\uparrow$10.95)
$\checkmark$ $\checkmark$ $\checkmark$ 95.40 98.43 86.94 80.03 90.20($\uparrow$15.85)
Table 3: Ablation study under the cross-dataset setup of the Consistency branch (C), Heatmap branch (H), and E-FPN.
@@
Table 4: Traditional FPN versus E-FPN, using the SBI-based data synthesis under the cross-dataset protocol. Bold and Underline indicate the best and the second-best performance, respectively. We report the results when integrating features $\mathbf{F}^{(i)}$ from different layers.
@@ 5 Conclusion
Acknowledgment
This work is supported by the Luxembourg National Research Fund, under the BRIDGES2021/IS/16353350/FaKeDeTeR and UNFAKE, ref.16763798 projects, and by POST Luxembourg.
@@ 8 Additional Results
Table 6 and Table 7 report the results under the in-dataset and the cross-dataset settings, respectively. Overall, it can be seen that LAA-Net achieves better performances than other state-of-the-art methods.
@@ 8.1 Qualitative Results: E-FPN versus FPN
8.2 Qualitative Results: Gaussian Noise
Figure 10: Detection of vulnerable points w/o and w/ Gaussian noise.
In Table 2 of the main manuscript, the performance of LAA-Net declined significantly when encountering Gaussian Noise perturbations. One possible reason is that the introduction of noise elevates the difficulty of detecting the vulnerable points. To confirm that, we report the inference of the heatmap before and after applying a Gaussian Noise on a facial image in Figure 10. As it can be observed, the detection of vulnerable points is highly impacted with the introduction of a Gaussian noise.
8.3 Robustness to Compression
To assess the robustness of LAA-Net to compression, we test LAA-Net on the c23 version of FF++, and the overall AUC is equal to $89.30$%.